A person weighs 61 N on the surface of the moon and 378 N on the surface of the earth. What is the acceleration of gravity on the surface of the moon? Round your answer to one decimal place.
Answer #1

Mass on earth = mass on moon

378/g_earth = 61/ g_moon

378/ 9.8 = 61/ g_moon

g_moon = 1.6 m/s^2

Comment in case any doubt please rate my answer....